OpenStudy (anonymous):

solve using elimination: 5x+y=0 15x-3y=9

OpenStudy (anonymous):

sorry wrong question

hero (hero):

What the?

OpenStudy (anonymous):

STUPID COMPUTER!!!!!

hero (hero):

Why do you come here posting non-sense?

OpenStudy (anonymous):

My computer keeps on posting what I said on the last problem I did

OpenStudy (anonymous):

hero do you know implicit differentiation?

hero (hero):

A little bit, but I already see what you're working on. I don't really feel like dealing with that right now.

OpenStudy (anonymous):

ok. There is just noone else on who knows calculus

hero (hero):

Talk to James. He eats Calculus for breakfast.

OpenStudy (anonymous):

sooooooo ummm about this problem.....

hero (hero):

oops

OpenStudy (anonymous):

Sorry for interrupting

hero (hero):

Write both equations in terms of y, then set y = y and solve for x.

OpenStudy (anonymous):

explain if u don't mind

hero (hero):

Is it big enough?

hero (hero):

By the way, that's wrong

OpenStudy (anonymous):

so x = -48/32?

OpenStudy (anonymous):

\[\huge\matrix{\text{5x+y=0}\\\text{15x-3y=9}}\] multiply top row by 3 and subtract \[\huge\matrix{\text{15x+3y=0}\\\text{15x-3y=9}\\6y=-9\Rightarrow y=-\frac{3}{2}}\] \[\huge 5x-\frac{3}{2}=0\Rightarrow x = \frac{3}{10}\]